# High Resolution Millimeter Wave Imaging Based on FMCW Radar Systems

### v3

@@ Abstract
In this paper, we present a unique 2D high resolution, compact, low-cost, low-weight, and highly accurate millimeter wave imagery system capable of operating in all weather conditions. We describe millimeter wave imaging process in detail and present several novel signal processing methods with their applications. To create the array, we utilize the Synthetic Aperture Radar (SAR) concept. The imagery system presented in this paper, can strongly compete with Lidar systems as the resolution limit is at the same level. Furthermore, in contrast to the Lidar systems, our imagery system can operate in heavy rain and dense fog and produce high quality images.
@@ I Introduction
Recently, millimeter, sub-terahertz, and terahertz waves have been used to perform radar imaging [11, 12, 13, 14, 15, 16, 17, 5, 6, 7, 8, 9, 10]. Compared to the systems that operate at lower frequencies, the advantage of the systems which operate at millimeter wave, sub-terahertz, and terahertz frequencies is that they are light-weight, compact, and have higher bandwidth. Operating with higher bandwidth will allow the system to obtain higher resolution in the range direction. In addition, operating at higher frequencies makes it possible to create apertures with large number of elements while the array still occupies a small physical space. This, in turn, yields higher resolution in the azimuth direction.
@@ I-B Motivation
In this paper, we present a high resolution imagery system. The system which we present in this paper, is our custom-made wide-band FMCW radar system. We have selected the W-band with frequencies ranging from 75 GHz to 108 GHz with 33 GHz bandwidth. Compared to the systems operating at lower frequencies, the W-band will make it possible to design a compact and light-weight system. In addition, we can reach higher bandwidth which as a result allows to enhance the range resolution limit of the system.
@@ II System Model
where $f_{c}$ is the carrier frequency and $\sigma_{l}$ is the radar cross section for the $l^{\text{th}}$ point target. The parameter $\beta$ is given as ${b/T}$, where $b$ and $T$ stand for the bandwidth and the chirp-time, respectively. Also, $u$ represents the envelope of the signal and is considered to be a rectangular window. Furthermore, $t$ and $\eta$ represent the fast- and slow-time parameters, respectively. Finally, $R_{l}$ is the radial distance between the target and the radar.
@@ III Apodization
[…]
@@
$\begin{cases}&\displaystyle\frac{\partial u}{\partial t}=\nabla.\left({c(|\nabla(G_{\sigma}*u)|^{2})}{\nabla u}\right),\\ &u(x,0)=f(x),\\ &c(||\nabla I||)=e^{-\displaystyle{\left(||\nabla I||/K\right)}^{2}},\\ &G_{\sigma}(I)=\sigma^{-1/2}e^{\displaystyle-|I|^{2}/(4\sigma)},\\ &(t,x)\in(0,T)\times\Omega.\end{cases}$ (5)
In (LABEL:denoising_opt_EL), the parameter K controls the sensitivity to the edges and is a hyper-parameter and ()(*)( ∗ ) represents the convolution operator. The algorithm presented in (LABEL:denoising_opt_EL) is, in fact, the nonlinear diffusion method, also known as modified Perona-Malik diffusion technique [24], which is used for image de-noising.
@@ VIII Conclusion
The paper is a proof that images with high resolution and high quality can be generated using compact, low-weight, and low-cost millimeter wave systems such as the one presented in this work.

### v4

@@ Abstract
In this paper, we present a unique 2D high resolution, compact, low-cost, lightweight, and highly accurate millimeter wave imagery system capable of operating in all weather conditions. We describe millimeter wave imaging process in detail and present several novel signal processing methods with their applications. To create the array, we utilize the Synthetic Aperture Radar (SAR) concept. The imagery system presented in this paper, can strongly compete with Lidar systems as the resolution limit is at the same level. Furthermore, in contrast to the Lidar systems, our imagery system can operate in heavy rain and dense fog and produce high quality images.
@@ I Introduction
Recently, millimeter, sub-terahertz, and terahertz waves have been used to perform radar imaging [11, 12, 13, 14, 15, 16, 17, 5, 6, 7, 8, 9, 10]. Compared to the systems that operate at lower frequencies, the advantage of the systems which operate at millimeter wave, sub-terahertz, and terahertz frequencies is that they are lightweight, compact, and have higher bandwidth. Operating with higher bandwidth will allow the system to obtain higher resolution in the range direction. In addition, operating at higher frequencies makes it possible to create apertures with large number of elements while the array still occupies a small physical space. This, in turn, yields higher resolution in the azimuth direction.
@@ I-B Motivation
In this paper, we present a high resolution imagery system. The system which we present in this paper, is our custom-made wide-band FMCW radar system. We have selected the W-band with frequencies ranging from 75 GHz to 108 GHz with 33 GHz bandwidth. Compared to the systems operating at lower frequencies, the W-band will make it possible to design a compact and lightweight system. In addition, we can reach higher bandwidth which as a result allows to enhance the range resolution limit of the system.
@@ II System Model
where $f_{c}$ is the carrier frequency and $\sigma_{l}$ is the radar cross section for the $l^{\text{th}}$ point target. The parameter $\beta$ is given as b/T{b/T}, where $b$ and $T$ stand for the bandwidth and the chirp-time, respectively. Also, $u$ represents the envelope of the signal and is considered to be a rectangular window. Furthermore, $t$ and $\eta$ represent the fast- and slow-time parameters, respectively. Finally, $R_{l}$ is the radial distance between the target and the radar.
@@ V Image De-noising
To perform the image de-noising process, we implement the $\rm l1$-TV method [21] which is based on the following convex optimization problem,
@@
[…]
In (LABEL:denoising_opt_EL), the parameter K controls the sensitivity to the edges and is a hyper-parameter and ()(*) represents the convolution operator. The algorithm presented in (LABEL:denoising_opt_EL) is, in fact, the nonlinear diffusion method, also known as modified Perona-Malik diffusion technique [24], which is used for image de-noising.
@@ VIII Conclusion
The paper is a proof that images with high resolution and high quality can be generated using compact, lightweight, and low-cost millimeter wave systems such as the one presented in this work.